OpenStudy (anonymous):

Find f^(xxvi)(x) if f (x) = sin x, So its asking for the 26th derivative.

OpenStudy (anonymous):

The first five derivatives is the following list: {cos(x),-sin(x),-cos(x),sin(x),cos(x)} Mod(26,4)=2 The 26th derivative is -sin[x]

OpenStudy (mr.math):

\(f(x)=\sin{x}\) \(f'(x)=\cos{x}\) \(f''(x)=-\sin{x}\) \(f^{(3)}=-\cos{x}\) \(f^{(4)}=\sin{x}\) From this pattern you can see that the derivatives repeat themselves after each four derivatives, Thus \(f^{(26)}=f^{(2)}=-\sin{x}\).

OpenStudy (anonymous):

Thanks. I've been getting these wrong for the stupid mistake that I started the sequence at sin x insted of cosine